# Applied-field magnetic structure and spectroscopy shifts of the effective spin-, -like magnet Li2CoCl4

### v1

@@ IV.2 Paramagnetic free-ion model
Also of note is the narrow absorption peak within the broader continuum of Fig. 7a at high fields. The line shape is well-fit by a single peak, indicating an isotropic -factor. Therefore, the peak may originate from a Co2+ site with minimal axial distortion. A similar peak was observed in the X-band EPR spectrum of LiCoO2 and was ascribed to a high-spin Co2+ site on the surface of polycrystalline sample [31, 32] . Powder spectrum simulations with an effective spin- model that consider a mixture of isotropic (surface) and anisotropic (bulk) Co2+ sites (Fig. 7b) indicate that 2% of the Co2+ sites are isotropic. We will regard the fraction as a small paramagnetic impurity and will ignore it while interpreting other magnetic property measurements of powder samples.
Still, for the extracted -factors, we need to justify an effective spin- model for high-spin (spin-) Co2+ in Li2CoCl4, with the only previous evidence of applicability being entropy changes extracted from the magnetic heat capacity [18]. Observed -factor values are well-studied for high-spin Co2+ with a local octahedral environment [33, 34, 35, 36]. Under a cubic field, the ground state of Co2+ is a (4F) orbital triplet with = 1 and = , and the following Hamiltonian describes the low-energy diagram [35]:
@@
is the spin-orbit coupling constant, is the crystal field energy due to axial distortion, B is the magnetic field vector, is the free-electron -factor, and is the orbital reduction factor. Without axial distortion or a magnetic field, the first-order spin-orbit coupling splits the twelve-fold degenerate state, making a = doublet the lowest energy level. The ground state doublet produces effective spin- behavior with an isotropic -factor of =(10+2)/3. The covalency of metal-ligand bonds or admixture of higher-energy orbital states, e.g. (4P), can produce values from 1 to 1.5, giving isotropic -factors between 4 and 4.33 [37].
@@
Axial distortion of the ideal octahedral environment of the Co2+ ion further splits the manifold, leading to a different value when a magnetic field is applied along or perpendicular to the distortion axis. For distortion values () small relative to the spin-orbit exchange constant (), the -factor anisotropy is approximately [37]:
@@
The average -factor in this case remains independent of the distortion effects since . For our data, is 4.41, which agrees well with an isotropic -value of 4.33 corresponding to high-spin Co2+ ion in the weak crystal field limit ( = 1.5). This spectroscopic determination of the high average -factor also corroborates the neutron diffraction moment, which we would expect to equal . Therefore, the observed -factors for Li2CoCl4 confirm a ground state, allowing for the modeling of the material as an effective spin- system at low temperatures.
@@
Solving the free Co2+ ion Hamiltonian (Eq. 1) further allows us to probe the interplay between the spin-orbit exchange constant and crystal field axial distortion in Li2CoCl4. Using the EasySpin package [28], we tuned the axial distortion and calculated the resulting -factor dependence, assuming an isotropic orbital reduction factor (Fig. 8a). The best match to the experimental -factor values is =2 with an adjusted parameter of 1.77, which is close to the weak crystal field case of 1.5. Such an increase in the reduction factor is reasonably similar to a previous analysis of cobalt salts [33] and demonstrates the deficiencies of neglecting the magnetic coupling between neighboring Co2+ ions [37].
@@
Besides its influence on the -factor, the octahedral distortion also substantially modifies the energy levels of the state, causing a splitting into six Kramers doublets. The excitation energy, i.e. the energy difference between the higher-energy Kramer doublets and the ground state, is calculated and shown in Fig. 8b in dimensionless units of the spin-orbit exchange constant. An applied magnetic field further splits the Kramers doublets, allowing us to distinguish the corresponding spectroscopic peaks from other spectroscopic features in the transmission spectrum. Indeed, a pronounced absorption peak at 644 cm-1 reveals a noticeable shift and broadening under applied magnetic fields (Fig. 9). We attribute this energy to a transition from the ground state to the second excited Kramers doublet, highlighted by a blue circle in Fig. 8b. Using the previously estimated values of and , this energy yields a spin-orbit coupling constant of = 157 cm-1, consistent with the 178 cm-1 value for a free Co2+ ion [33]. The transition to the first excited state, however, is not observed in the transmission spectrum because of a strong absorption band between 200 and 400 cm-1. Nevertheless, additional field-dependent spectral features (blue dots in Fig. 8b) were observed at higher energies (Fig. S8 [29]) and are likely associated with the remaining inter-Kramers doublet transitions. While the free-ion model does not provide a complete quantitative description of the IR transmission spectrum, it qualitatively approximates the underlying physics of the Co2+ ion in Li2CoCl4.
IV.3 Spin gaps below the saturation field
Within the free-ion model framework, the observed 11.6 cm-1 gap could be interpreted as the zero-field energy of a system. Exploring this possibility, we found moderate agreement between the experimental absorption and a Hamiltonian with a single-ion anisotropy of cm-1 and anisotropic -factors of and (Fig. S9 [29]). However, to observe such a low-energy first excited Kramers doublet would require extremely high axial distortions (Fig. 8b), which would then require -factors far from those obtained in simulations (Fig. 8a). Unsurprisingly, therefore, understanding the low-field data requires a model involving spin-spin exchange interactions.
The 11.6 cm-1 and 17.2 cm-1 spin gaps also cannot be explained using the phenomenological theory of antiferromagnetic resonance (AFMR) for collinear magnets with axial magnetic anisotropy [38]. The model predicts two excitation modes with spin-flop and spin-flip reorientations with increasing magnetic field. The magnetic field softens the AFMR resonance energies until the spin gaps close at the critical fields, which are directly related to the zero-field energies by the -factor. But when we use the experimental spectroscopic -factors, the calculated energies are significantly lower than the observed spin gaps, meaning the critical fields would need to be much higher than those observed in the magnetization data for the spin gaps to be AFMR modes. Again, a more involved spin-spin interaction model is needed to understand the spectroscopic data in the lower field regions.
IV.4 Paramagnetic magnetization and susceptibility
@@
For Li2CoCl4, spin-orbit coupling is critical to understanding the susceptibility and magnetization, which are both influenced by the material’s low-energy Kramers doublets. Fig. 10 compares the experimental data in the paramagnetic region to EasySpin calculations utilizing Eq. 1 for three cases: a spin- model with the spectroscopic anisotropic -factors, a spin-orbit coupling model with the experimental spectroscopic parameters, and a spin-orbit coupling model with optimized spectroscopic parameters. All three models qualitatively match the shape of the magnetization data and indicate a saturated moment near 2–2.75 , but they differ significantly in describing the susceptibility data. While the spin-orbit coupling models reasonably match the low- and high-temperature slopes of the susceptibility, the spin- model only captures the low-temperature behavior. Improvements to the spin-orbit coupling model would need to consider higher-order spin-orbit interactions and electronic structure effects.
IV.5 Evidence of -like interactions
@@
Our calculations (Fig. 8b) estimate an energy gap between the two lowest doublets of 264 cm-1 (32.7 meV), indicating that Co2+ ions in Li2CoCl4 remain in the ground state Kramers doublet at low temperatures. This state can be mapped to an effective spin- state with an exchange interaction, , between next-neighbor spins exhibiting anisotropy of the type . Solving the spin-orbit Hamiltonian (Eq. 1) using spectroscopically determined parameters enables an estimation of the magnetic coupling anisotropy, yielding for Li2CoCl4 (see the supplemental material for details [29]). This value of classifies Li2CoCl4 as a system with easy-plane anisotropy, mirroring the -factor anisotropy. Most Co-based chain materials, such as CsCoCl3 [19], CoCl2H2O [39], CoNb2O6 [40], [Co(NCS)2(4-methoxypyridine)2]n [41], Co(N2H5)2(SO4)2 [42], and (Ba,Sr)Co2V2O8 [43], exhibit Ising anisotropy (). We believe that the tetragonal distortion of the CoCl6 octahedron in Li2CoCl4, in contrast to the trigonal distortion in those systems, may serve as the underlying mechanism for anisotropy.
The neutron diffraction data is also compatible with -plane anisotropy. Though the magnetic intensity changes in the polycrystalline data support a re-orientation along , there were no observed reflections that violate a rotation through the axis, and the refinement with moment components along and (2 space group) was nearly identical to that with moments only along . In contrast, the noted increase of the (020) reflection indicates that the magnetic moments are not rotating through the axis. Previous DFT results using an experimental unit cell also indicate that moments along or are extremely similar energetically [18].
@@
Qualitatively, the magnetic heat capacity of Li2CoCl4 at temperatures just above the long-range magnetic ordering [18] displays a hump similar to that of the predicted heat capacity for an isolated spin-, chain with , , and an applied field along the direction [44]. Fitting the heat capacity of Cs2CoCl4 [45], a compound with tetrahedrally-coordinated Co2+ which also has a long-range ordering lambda peak and a hump in its magnetic heat capacity, to this model gives an intrachain exchange interaction parameter. However, we could not similarly fit the heat capacity of polycrystalline Li2CoCl4, which has appreciable interchain interactions. Our neutron diffraction data suggest that the hump in the magnetic heat capacity stems from short range order due to the spin-flop transition.
V Conclusions
@@
Li2CoCl4 exhibits strongly anisotropic magnetic properties. Neutron diffraction suggests that the moments within the octahedrally coordinated chains of Co2+ prefer alignment along the crystallographic axis, though an -plane canted moment could not be ruled out under applied magnetic fields. At zero field, the moments align ferromagnetically within the chains and antiferromagnetically between nearest-neighboring chains. At nonzero fields below 1.6 T, neutron diffraction shows that the antiferromagnetic structure is unchanged, but with increasing magnetic field, the anti-aligned chains undergo a spin-flop transition to a field-aligned ferromagnetic state, likely with symmetry. For the randomly oriented crystallites in a polycrystalline sample, a majority of the chains are aligned near 3.5 T. The chains do not transition to a ferrimagnetic state at intermediate fields as in CoCl2H2O.
@@
Anisotropic -factors and magnetic spectroscopy data confirm a well-separated Kramers doublet ground state ( = ) for the Co2+ ion that is 32.7 meV below the first excited state. Consequently, Li2CoCl4 can be modeled as an effective spin- system of Co2+ chains with significant spin-orbit coupling and interchain interactions. The = ground state and strong spin-orbit coupling explain the qualitative shape of magnetization and susceptibility data in the paramagnetic region, as well as the refined magnetic moment from neutron diffraction, the entropy recovery extracted from previous heat capacity data, and the difficulty in fitting the paramagnetic susceptibility to a purely Curie-Weiss shape. The tetragonal distortion of the CoCl6 octahedra may also produce -like magnetic interactions, with a calculation of the interaction parameter ratio / yielding the value 0.41. anisotropy distinguishes Li2CoCl4 from other Co2+ chain magnets that have trigonally distorted octahedra and display Ising-like anisotropy.
@@
Further work should explore this anisotropy and its influence on the unexplained spin gaps below the moment saturation field by developing a model of spin-spin interactions between the cobalt ions. The models will be aided by single crystal neutron scattering experiments. Single crystal magnetization and neutron data along the three crystallographic axes would also improve the description of the magnetic behavior, allowing for a more precise determination of the moment saturation field and for an evaluation of potential magnetic moment canting within the plane. Overall, Li2CoCl4 offers a rare opportunity to study the combined effects of effective spin- behavior and -like magnetism on spin dynamics in an octahedrally coordinated Co2+ system.

### v3

@@ III.3.1 Noninteracting spin- model
Also of note is the narrow feature within the broader continuum of Fig. 5a at high fields. The lineshape is well-fit by a single peak, indicating an isotropic -factor of 4.19. The peak may originate from a Co2+ site with minimal axial distortion. A similar peak was observed in the X-band EPR spectrum of LiCoO2 and was ascribed to a high-spin Co2+ site on the surface of polycrystalline sample [36, 37]. Powder spectrum simulations that consider a mixture of isotropic (surface) and anisotropic (bulk) Co2+ sites (Fig. 5b) indicate that 2% of the Co2+ sites are isotropic. We will regard the fraction as a small paramagnetic impurity and will ignore it while interpreting other magnetic property measurements of powder samples. The fit -factors in the noninteracting spin- model are significantly larger than the isotropic, free-electron value of . The discrepancy can be well explained by an effective Hamiltonian that accounts for spin-orbit coupling.
III.3.2 Effective spin-orbit coupling model
To build a spin-orbit coupling model, we first consider the ligand environment and paramagnetic susceptibility of Co2+ in Li2CoCl4. The Co2+ cations are octahedrally coordinated to Cl- anions, and the magnetic susceptibility indicates that the Co2+ cations are high-spin with unquenched orbital angular momentum [18], i.e. significantly influenced by spin-orbit coupling (SOC). The anisotropic magnetism of a high-spin Co2+ cation in a local octahedral environment is well-studied [38, 39, 40, 41, 32]. The ground state is a (4F) orbital triplet with = 1 and = , and the following Hamiltonian describes the low-energy diagram [41]:
@@ III.3.2 Effective spin-orbit coupling model
is the spin-orbit coupling constant, is the crystal field energy due to axial distortion, B is the magnetic field vector, is the free-electron -factor, and is the orbital reduction factor. The orbital reduction factor quantifies the covalence of metal-ligand bonds and admixture of higher-energy orbital states, e.g. (4P). Without axial distortion or a magnetic field, the first-order spin-orbit coupling splits the twelve-fold degenerate state, making a = doublet the lowest energy level. The doublet is separated by from the first excited quartet state (49.7 meV assuming a free-ion spin-orbit coupling constant of 178 cm-1) [38, 42, 40]. The well-separated, ground-state doublet, therefore, produces effective spin- behavior at low temperatures with an isotropic -factor of =(10+2)/3 with values ranging from 1 (strong crystal field) to 1.5 (weak crystal field), resulting in isotropic -factors between 4 and 4.33 [43].
@@
Next, axial distortion should be considered because of the aforementioned tetragonal distortion of the CoCl6 octahedra in Li2CoCl4. Axial distortion of the ideal octahedral environment of the Co2+ ion further splits the manifold into six Kramers doublets and leads to anisotropy of the -factor. For distortion values () small relative to the spin-orbit exchange constant (), the -factor anisotropy is approximately [43]:
@@
The average -factor in this case remains independent of the distortion effects since . Using the fit spectroscopic -factors and , we can calculate a of 4.41, which agrees well with the predicted isotropic -factor of 4.33 for high-spin Co2+ ion in the weak crystal field limit. The calculated also corroborates the neutron diffraction moment, which we would expect to equal if . Therefore, the observed -factors for Li2CoCl4 confirm a (Kramers doublet) ground state.
@@
Having confirmed the applicability of the spin-orbit coupling model, we can solve the Hamiltonian (Eq. 1) to further probe the interplay between the spin-orbit exchange constant and crystal field axial distortion in Li2CoCl4. Using the EasySpin package [28], we tuned the ratio of the axial distortion to the spin-orbit exchange constant, /(), and calculated the resulting changes in and (Fig. 6a). The experimental -factor values are well-described by a ratio of 2 and a reduction factor () of 1.77, which is close to the weak crystal field case of 1.5. The relatively small ratio [38] is indicative of a system where spin-orbit coupling plays a significant role [41]. The reduction factor is reasonably similar to that of other cobalt salts [38], though the value is outside the expected limits of [1,1.5] for a noninteracting system, highlighting a deficiency of neglecting the magnetic coupling between neighboring Co2+ ions [43].
III.3.3 Observed excitations predicted by the SOC model
@@
To assess the SOC model at higher energies in the absorption spectrum, we calculated the energy difference between the ground-state and higher-energy Kramers doublets as a function of the ratio (Fig. 6b) to compare with the experimental data. Since an applied magnetic field further splits the Kramers doublets, we can distinguish spectroscopic peaks associated with the transitions from other spectroscopic features in the experimental spectrum. For example, a pronounced peak at 644 cm-1 shifts and broadens under an applied magnetic field (Fig. 7). Assuming , we attribute this energy to a transition from the ground state to the second excited Kramers doublet, denoted by a blue circle in Fig. 6b. Using the previously estimated value of (1.77), this energy yields a spin-orbit coupling constant of = 157 cm-1, consistent with the 178 cm-1 value for a free Co2+ ion [38]. The transition to the first excited state, however, is not observed in the spectrum because of a strong absorption band between 200 and 400 cm-1. Nevertheless, additional field-dependent spectral features (blue dots in Fig. 6b) were observed at higher energies (Fig. S9 [29]) and are likely associated with the remaining inter-Kramers doublet transitions. The model, therefore, qualitatively captures the anisotropic physics of Co2+ in Li2CoCl4 at higher energies in the measured absorption spectrum.
III.3.4 Paramagnetic magnetization and magnetic susceptibility
@@
In addition to predicting IR excitations, the SOC model can be compared to experimental magnetization and magnetic susceptibility data, where the contributions from low-lying excited states are relevant. Fig. 8 compares the experimental data in the paramagnetic region to EasySpin calculations utilizing Eq. 1 for three cases: a noninteracting spin- model with the -factors from our high-field spectroscopy data; a spin-orbit coupling model with the , , and parameters fit to our high-field spectroscopy data; and a spin-orbit coupling model with spectroscopic parameters optimized to fit the experimental magnetization and magnetic susceptibility. All three models qualitatively match the shape of the magnetization data and indicate a saturated moment near 2-2.75 , but they differ significantly in describing the susceptibility data. While both spin-orbit coupling models reasonably match the low- and high-temperature slopes of the susceptibility, the spin- model only captures the low-temperature behavior. Improvements to the spin-orbit coupling model would need to consider higher-order spin-orbit interactions and electronic structure effects.
III.3.5 Spin gaps below the saturation field
The approximation of free spin dynamics in high magnetic fields, of course, breaks down at lower magnetic fields where the spin-spin interactions are dominant in determining the excitation spectrum. At zero field, we observe a strong peak at 11.6 cm-1 and a weaker satellite at 17.2 cm-1 (Fig. 5c) that we attribute to spin gaps in the collective spin excitations of Li2CoCl4. Within a free-ion framework, the observed 11.6 cm-1 gap could be interpreted as the zero-field energy of a system. Exploring this possibility, we found moderate agreement between the experimental data and a Hamiltonian with a single-ion anisotropy of cm-1 and anisotropic -factors of and (Fig. S10 [29]). However, to observe a first excited Kramers doublet 10 cm-1 would require extremely high axial distortions (Fig. 6b), which would be inconsistent with the model’s -factors (Fig. 6a).
The 11.6 cm-1 and 17.2 cm-1 spin gaps also cannot be explained using the phenomenological theory of antiferromagnetic resonance (AFMR) for collinear magnets with two-axis magnetic anisotropy [44]. The model predicts two excitation modes attributed to spin-flop and spin-flip reorientations with increasing magnetic field. The magnetic field softens the AFMR resonance energies until the spin gaps close at the critical fields, which are directly related to the zero-field energies by the -factor. But for a spin-flop field of 1.6 T () and a resonance frequency of 11.6 cm-1, the necessary -factor would be , significantly above our observed -factors. For the features to be AFMR modes, the zero-field gap would have to be much smaller or the spin-flop field much larger. Therefore, understanding the low-field spectroscopic data requires a more detailed microscopic model involving anisotropic spin-spin exchange interactions.
III.4 Evidence of -like anisotropy
@@
A notable implication of our IR spectroscopy data is -like magnetism of the Co2+ cation in Li2CoCl4 since . Transition energies calculated with the SOC model (Fig. 6b) predict an energy gap between the two lowest doublets of 264 cm-1 (32.7 meV), indicating that Co2+ ions’ behavior in Li2CoCl4 is dominated by the ground-state Kramers doublet at low temperatures. This state can be mapped to an effective spin- state with an exchange interaction, , between next-neighbor spins exhibiting anisotropy of the type . Solving the spin-orbit Hamiltonian (Eq. 1) using spectroscopically determined parameters enables an estimation of the magnetic coupling anisotropy, yielding for Li2CoCl4 (see the supplemental material for details [29] and references [38, 31, 19, 32] therein). This value of classifies Li2CoCl4 as a system with easy-plane anisotropy, consistent with . Most Co-based chain materials with octahedral Co coordination, such as CsCoCl3 [19], CoCl2H2O [45], CoNb2O6 [33], [Co(NCS)2(4-methoxypyridine)2]n [46], Co(N2H5)2(SO4)2 [47], and (Ba,Sr)Co2V2O8 [34], exhibit Ising anisotropy (). The tetragonal distortion of the CoCl6 octahedron in Li2CoCl4, in contrast to the trigonal distortion in those systems, may serve as the underlying mechanism for anisotropy.
The neutron diffraction data are also compatible with, though not definitive regarding, easy-(-)plane anisotropy. The data suggest a preference for moment alignment along the axis, but there were no observed reflections that violate a rotation through the axis during the spin reorientation, and the refinement with moment components along and (2 space group) was nearly identical to that with moments only along . In contrast, the zero-field structure has no moment component along , and under an applied field, the maximal magnetic space group with moments solely along was incompatible with the magnetic intensity. These features of the neutron data suggest that the axis or plane may be the material’s easy axis/plane, which is consistent with the IR data. If the system has easy-axis anisotropy along instead of easy-plane anisotropy, this is still compatible with the -anisotropy description, which neglects spin-spin exchange interactions that would, presumably, distinguish the and axes. The SOC model was also overparameterized if we introduced additional anisotropy to try to distinguish and , leaving anisotropy as the best description with the available data.
@@
Qualitatively, the magnetic heat capacity of Li2CoCl4 at temperatures just above the long-range magnetic ordering [18] displays a hump similar to that of the predicted heat capacity for an isolated spin-, chain with , , and an applied field along the direction [48]. Fitting the heat capacity of Cs2CoCl4 [49], a compound with tetrahedrally-coordinated Co2+ which also has a long-range ordering lambda peak and a hump in its magnetic heat capacity, to this model gives an intrachain exchange interaction parameter. However, we could not similarly fit the heat capacity of polycrystalline Li2CoCl4, which has appreciable interchain interactions. The hump is, therefore, better understood as stemming from short-range order due to the spatially distributed spin reorientation with increasing field.
IV Conclusions
@@
Overall, Li2CoCl4 exhibits strongly anisotropic magnetic properties. At zero field, neutron diffraction shows that the moments within chains of octahedrally coordinated Co2+ align ferromagnetically while aligning antiferromagnetically between nearest-neighboring chains. At nonzero fields below 1.6 T, neutron diffraction shows that the antiferromagnetic structure is unchanged, but with increasing magnetic field, the anti-aligned chains undergo a spin reorientation to a ferromagnetic state, likely with a preference for moment alignment along the axis. Based on refinements including the antiferromagnetic phase and a ferromagnetic phase, a majority of the sample becomes ferromagnetic near 3.5 T. The chains do not transition to a ferrimagnetic state at intermediate fields as in CoCl2H2O.
@@
The high-field FIR data are well-described by a spin-orbit coupling model that confirms a well-separated Kramers doublet ground state ( = ) for the Co2+ ion that is approximately 32.7 meV below the first excited state. The = ground state and strong spin-orbit coupling explain the qualitative shape of magnetization and susceptibility data in the paramagnetic region, as well as the refined magnetic moment from neutron diffraction, the entropy recovery extracted from previous heat capacity data, and the difficulty in fitting the paramagnetic susceptibility to a purely Curie-Weiss shape. Still, the two spin gaps at 11.6 and 17.2 cm-1 in the zero-field spectrum highlight the need for a model that also accounts for the anisotropy of spin-spin interactions. The tetragonal distortion of the CoCl6 octahedra may also promote -like magnetic interactions, with an estimated interaction parameter ratio / of 0.41. anisotropy distinguishes Li2CoCl4 from other Co2+ chain magnets that have trigonally distorted octahedra and display Ising-like anisotropy.
@@
Further work should explore this anisotropy and its influence on the unexplained spin gaps below the moment saturation field by developing a model of spin-spin interactions between the cobalt ions. The models will be aided by single crystal neutron scattering experiments. Single crystal magnetization and neutron data along the three crystallographic axes would also improve the description of the magnetic behavior, allowing for a more precise determination of the moment saturation field and a definitive description of the spin reorientation at intermediate magnetic fields. Li2CoCl4, therefore, offers a rare opportunity to study the combined effects of effective spin- behavior and -like magnetic anisotropy on spin dynamics in an octahedrally coordinated Co2+ system.
@@ Acknowledgements.
- Weil and Bolton [2007] J. A. Weil and J. R. Bolton, Electron Paramagnetic Resonance: Elementary Theory and Practical Applications (Wiley-Interscience, Hoboken, NJ, 2007) pp. 99–105.
@@
- Mabbs and Machin [1973] F. E. Mabbs and D. J. Machin, Magnetism and Transition Metal Complexes (Chapman and Hall, London, 1973).
@@
- Goodenough [1968] J. B. Goodenough, Spin-orbit-coupling effects in transition-metal compounds, Physical Review 171, 466 (1968).